OpenStudy (anonymous):

Use the discriminant to determine the number of real-number solutions for the equation: 8x2 + 8x + 2 = 0

OpenStudy (anonymous):

. A.one solution B. two solutions C. no solutions D. infinitely many solutions

OpenStudy (anonymous):

i think the aswer is d

OpenStudy (lυἶცἶ0210):

What did you get for the discriminant?

OpenStudy (anonymous):

discriminant=> b^2-4*a*c=0

OpenStudy (anonymous):

the given equation is of the form:- ax^2+bx+c=0

OpenStudy (anonymous):

can u tell me the values of a,b and c from the given equation? can u identify it?

OpenStudy (anonymous):

yhea well i have to find out if the equation is one solution two solutions no solutions infinitely many solutions

OpenStudy (triciaal):

general form for the quadratic equation ax^2 + bx + c = 0 discriminant, D = b^2 - 4ac when D = 0 1 real root when D < 0 no real roots wen D > 0 2 real roots

OpenStudy (anonymous):

okay